how do you find the values for m and b in the following systems so that the solutions to the system is (-3,4)
5x+ 7y=b and mx+y=22

Answer by Earlsdon(6103)

Simply substitute the given values of x and y (-3, 4) into the two equations and solve for the respective values of b and m.
1)
5x + 7y = b Substitute x = -3 and y = 4.
5(-3) + 7(4) = b Simplify.
-15 + 28 = b
b = 13
2)
mx + y = 22 Substitute x = -3 and y = 4.
m(-3) + 4 = 22 Subtract 4 from both sides of the equation.
-3m = 18 Finally, divide both sides of the equation by -3.
m = -6